OpenStudy (aye_my_peeps):

WHAT IS 0.28 WRITTEN AS A FRACTION IN SIMPLEST FORM?

OpenStudy (vhtran):

0.28 in a fraction not simplified is 28/100

OpenStudy (vhtran):

If you go slow, step by step, it is easier to simplify by 2

OpenStudy (vhtran):

28 divided by 2 is 14, 100 divided by 2 is 50. So you get 14/50, repeat again. 7/25

OpenStudy (vhtran):

*sorry I meant divided by not simplify

OpenStudy (aye_my_peeps):

CHIOCES: A. \[\frac{ 1 }{ 5 } B. \frac{ 1 }{ 4 } C. \frac{ 7 }{ 25 } D. 2\frac{ 4 }{ 5 }\]

OpenStudy (vhtran):

C

OpenStudy (aye_my_peeps):

SRRY AND THNK U

OpenStudy (vhtran):

Np :D